OpenStudy (anonymous):

Can you help me solve this problem,and Explain it to me (: -2/5x - 9 < 9/10

OpenStudy (ajprincess):

Is this ur question? \[-\frac{ 2 }{ 5 }x-9<\frac{ 9 }{ 10}\]

OpenStudy (anonymous):

Yes :)

OpenStudy (jiteshmeghwal9):

\[\frac{-2(x)}{5}-9<\frac{9}{10}\]

OpenStudy (jiteshmeghwal9):

\[{-2x-45\over5}<\frac{9}{10}\]

OpenStudy (jiteshmeghwal9):

\[10(-2x-45)<9(5)\]

OpenStudy (jiteshmeghwal9):

\[-20x-450<45\]

OpenStudy (jiteshmeghwal9):

Now add 450 both sides & then divide -20 both sides :)

OpenStudy (ajprincess):

Nw to solve this inequality u hav to isolate x. To do that plz perform the following steps. 1.Add 9 to both sides. 2.Multiply both sides by 5. 3.Divide by -2 both sides. When u divide by a negative number the inequality sign flips. see if u can do it:)

OpenStudy (jiteshmeghwal9):

The last point of @ajprincess is very correct in inequality if we divide a negative number the inequality sign flips :)

OpenStudy (anonymous):

x > -99/4 ?

OpenStudy (ajprincess):

yup perfect:)

OpenStudy (anonymous):

:D yay! Thanks :)

OpenStudy (ajprincess):

welcome:) nd gud work:)